According to the report “Population Distribution and Change: 2000 to 2010” from the U.S. Census Bureau, the U.S. population increased by 27.3 million people from 2000 to 2010. This was a 9.7% increase. According to these figures, what was the U.S. population in the years 2000 and 2010?
Year 2000:
Year 2010:

Answer by Theo(13342)

the united states population increased from 27.3 million people between 2000 and 2010.

that was a 9.7% increase.

let x = the population in 2000.

the increase is therefore .097 * x.

since the increase is 27.3 million people, then .097 * x = 27.3 million.

solve for x to get x = 27.3 / .097 = 281.443299 million.

the population in 2000 was 281.443299 million.

the increase between 2000 and 2010 was 27.3 million.

that means that the population in 2010 was 308.743299 million.

308.743299 / 281.443299 = 1.097.

that's an increase of (1.097 - 1) * 100 = 9.7%.

the solution looks good.

the solution is that there were 281.443299 million people in the united states in the year 2000 and 308.743299 million people in the united states in the year 2010.